# thanks: Corresponding authorthanks: Corresponding author

### v2

@@ IV Null geodesic equations and black hole shadow
To determine the shape of the shadow cast by a rotating Sen BH, we initially examine the null geodesics in this spacetime geometry. Utilizing the Hamilton-Jacobi equation, we analyze the null geodesic structure of the rotating Sen BH, as outlined below [57]
@@
with, $C$ is the Carter constant [57].
@@
where $r_{0}$ is the distance between the observer and the BH, and $\theta_{0}$ is the inclination angle between the observer’s line of sight and the rotational axis of the BH. In the particular case where the observer is situated on the black hole’s equatorial plane with $\theta_{0}=\pi/2$, we obtain:
@@
Figure. 7 illustrates the shadows of rotating Sen black holes for various selected values of the spin parameter $a$ and the parameter $q_{m}$. The black solid line represents the boundary of the shadow cast by a Kerr BH ($q_{m}=0$). The rotating Sen black holes shadows are not perfect circles. It is found that, for a fixed spin parameter $a$, the size of the BH shadow diminishes with increasing $q_{m}$. When $a=0.99$, the shape of the BH shadow appears D-shaped.
Figure 7: The shadows cast by rotating Sen black holes vary with the parameters $a$ and $q_{m}$.
Let us now shift our focus to the actual size and distortion of the shadow cast by a rotating Sen BH. We introduce two observables that serve to characterize the BH shadow: $R_{s}$ and $\delta_{s}$. Specifically, the parameter $R_{s}$ denotes the actual size of the shadow, whereas $\delta_{s}$ represents the distortion in the shadow’s shape. These parameters will allow for a more detailed analysis of the black hole’s shadow within the context of its observable characteristics. To determine the radius $R_{s}$, we use three points on the shadow boundary: the top point ($x_{t},y_{t}$), the bottom point ($x_{b},y_{b}$), and the rightmost point ($x_{r},0$). As defined in [58]:
@@
Here, $D$ represents the distance between the leftmost point ($x_{l},0$) of the shadow and the rightmost point ($\tilde{x}_{l},0$) of the reference circle. The variations in $R_{s}$ and $\delta_{s}$ for the shadows are depicted in Fig. 8. The parameter $q_{m}$ significantly affects both the apparent shape and size of the BH shadow. As illustrated in Fig. 8, an increase in $q_{m}$ leads to a gradual decrease in the shadow’s size and a corresponding increase in its distortion.
Figure 8: The variations of $R_{s}$ and $\delta_{s}$ with respect to $q_{m}$ for rotating Sen black holes are shown. In the left panel, the spin parameter $a=0.3$, while in the right panel, $a=0.5$.
V CONSTRAINTS FROM THE EHT OBSERVATIONS OF SUPERMASSIVE BLACK HOLES OF M87 AND SGR A
The EHT observations of M87 and Sgr A impose significant constraints on theoretical models of black holes. Through imaging the shadows and surrounding emissions of these supermassive black holes, the EHT data allows researchers to refine parameter constraints. Such observations yield unprecedented insights into the characteristics of black holes and their surroundings. In this section, we employ EHT observational data related to the supermassive black holes at M87 and Sgr A to constrain rotating Sen black holes.
V.1 Constraints from the EHT results of M87
@@
here, $A$ is the BH shadow area, and $d$ represents the distance between the BH and Earth, which is $d=16.8$ Mpc for M87. Figure. 9 shows the variation in the shadow angular diameter of rotating Sen black holes with respect to parameters $a$ and $q_{m}$, for viewing angles $\theta_{0}=17^{\circ}$ and $\theta_{0}=90^{\circ}$. The M87 shadow’s angular diameter, constrained within the $1\sigma$ bound of $39\mu as\leq\theta_{d}\leq 45\mu as$, imposes limits on the spin parameter $a$ and the parameter $q_{m}$ based on the inclination angle $\theta_{0}$. Specifically, for $\theta_{0}=17^{\circ}$, we find $0<a<0.529574M$ and $0<q_{m}<0.215561M$. When $\theta_{0}=90^{\circ}$, the constraints are $0<a<0.692145M$ and $0<q_{m}<0.215561M$.
[…]
Figure 9: Density plots of the shadow angular diameter $\theta_{d}$ for rotating Sen black holes are displayed. The left plot is for $\theta_{0}=17^{\circ}$, and the right one corresponds to $\theta_{0}=90^{\circ}$. The observed angular diameter of M87 is $\theta_{d}=42\pm 3\mu as$ within $1\sigma$. The black solid line denotes $\theta_{d}=39\mu as$.
V.2 Constraints from the EHT results of Sgr A
@@
Figure 10: Density plots of the shadow angular diameter $\theta_{d}$ for rotating Sen black holes at two viewing angles: $\theta_{0}=50^{\circ}$ (left) and $\theta_{0}=90^{\circ}$ (right). The solid black line indicates $\theta_{d}=50\mu as$, representing the maximum observed value for Sgr A, while the dashed black line denotes $\theta_{d}=46.9\mu as$, the minimum observed value.
@@ VI Conclusion
In this study, we have derived a rotating Sen BH solution by applying the NJA to the spherical symmetric Sen BH. We have further investigated the physical properties of the event horizon and ergosurface associated with rotating Sen black holes. Our analysis reveals that the number of horizons is determined by both the effective charge $q_{m}$ and the spin parameter $a$. Notably, while the shape of the outer ergosphere is sensitive to variations in $a$, the size of the outer horizon decreases monotonically with increasing $a$. Moreover, for a fixed value of the spin parameter, both the radii of the outer horizon and outer ergosphere shrink as $q_{m}$ increases.
Additionally, the Novikov-Thorne model was employed to analyze the properties of thin accretion disks around rotating Sen black holes. Under specific values of the spin parameter $a$, the influence of $q_{m}$ on the energy flux, radiation temperature, and luminosity spectrum of the thin disk has been systematically explored. The results show that the energy flux, radiation temperature, and luminosity spectrum exhibit similar behavior: they initially increase with radius, reach a peak, and then decrease. For a fixed value of the spin parameter $a$, all three quantities grow as the effective charge $q_{m}$ increases. Moreover, when $q_{m}$ is held constant, the peak values of the energy flux, radiation temperature, and luminosity spectrum diminish with increasing $a$.
Furthermore, we have analyzed how the parameters $a$ and $q_{m}$ influence both the shape and size of the shadow cast by rotating Sen black holes. The results demonstrate that, for a fixed spin parameter $a$, the shadow size decreases as $q_{m}$ increases. Through comparison with EHT observational data for M87 and Sgr A, we have determined constraints on the parameters of rotating Sen black holes across various inclination angles. The derived bounds indicate that, under the combined observational constraints, the effective charge $q_{m}$ must lie within the interval $0.133158M<q_{m}<0.215561M$. These results confirm that the rotating Sen BH model is consistent with the current high-resolution BH imaging observations, and provide a theoretical basis for further exploring the string-theoretic effects in extreme gravitational fields.

### v3

@@ IV Null geodesic equations and black hole shadow
To determine the shadow cast by a rotating charged BH with scalar hair, we study the null geodesics in this spacetime using the Hamilton-Jacobi formalism, following Carter [64]
@@
with $C$ being the Carter constant [64].
@@
where $r_{0}$ is the distance between the observer and the BH, and $\theta_{0}$ is the inclination angle between the observer’s line of sight and the rotational axis of the BH. In the particular case where the observer is situated on the BH’s equatorial plane with $\theta_{0}=\pi/2$, we obtain:
@@
Figure 5: Shadows of a rotating charged black hole with scalar hair are presented for varying spin parameter $a$ and scalar hair parameter $s$, with $Q=0.3$ fixed.
Figure 6: Shadows of a rotating charged black hole with scalar hair are shown for varying spin parameter $a$ and charge parameter $Q$, with $s=0.1$ fixed.
Figures 5 and 6 show two-dimensional profiles of the shadow cast by a rotating charged BH with scalar hair. Figure 5 focuses on the effect of the scalar hair parameter $s$ on the shadow, with the charge parameter fixed at $Q=0.3$. For fixed spin $a$, increasing $s$ from zero leads to a monotonic decrease in the apparent size of the shadow (i.e., the shadow boundary contracts toward the center). At high spin values (e.g., $a=0.9$), larger $s$ not only shrinks the shadow but also induces a distinctive cuspy edge, markedly different from the nearly circular shadows observed at low spin (e.g., $a=0.3$). A cuspy edge also emerges at $a=0.7$ as $s$ increases, but the effect is more pronounced for $a=0.9$. In contrast, Figure 6 investigates the influence of the charge parameter $Q$, with the scalar hair parameter fixed at $s=0.1$. Here, for fixed $a$, increasing $Q$ similarly reduces the shadow size monotonically, and the deviation from the uncharged case becomes increasingly evident. At high spins (e.g., $a=0.9$), larger $Q$ enhances the non-circular distortion and also gives rise to a cuspy edge. A common trend emerges from both figures: regardless of whether $Q$ or $s$ is held fixed, increasing the other parameter (i.e., $s$ or $Q$) reduces the shadow size for a given $a$, and cuspy edges are more readily formed at high spin. However, a key difference exists: for the same $a$, the scalar hair parameter $s$ has a stronger effect on shadow deformation than the charge parameter $Q$. For instance, at $a=0.7$, a cuspy edge appears when $s=0.42$, whereas no such feature is observed for $Q=0.6$, indicating that scalar hair induces more significant shape distortions than electric charge.
Let us now shift our focus to the actual size and distortion of the shadow cast by a rotating charged BH with scalar hair. We introduce two observables that serve to characterize the BH shadow: $R_{s}$ and $\delta_{s}$. Specifically, the parameter $R_{s}$ denotes the actual size of the shadow, whereas $\delta_{s}$ represents the distortion in the shadow’s shape. These parameters will allow for a more detailed analysis of the BH’s shadow within the context of its observable characteristics. To determine the radius $R_{s}$, we use three points on the shadow boundary: the top point ($x_{t},y_{t}$), the bottom point ($x_{b},y_{b}$), and the rightmost point ($x_{r},0$). As defined in [65]:
@@
Figure 7: Variations of the shadow radius $R_{s}$ (top row) and deformation $\delta_{s}$ (bottom row) with scalar hair parameter $s$ (left column, $Q=0.3$) and charge $Q$ (right column, $s=0.1$), for spin parameters $a=0.3$ (green) and $a=0.5$ (red).
Here, $D$ represents the distance between the leftmost point ($x_{l},0$) of the shadow and the rightmost point ($\tilde{x}_{l},0$) of the reference circle. Figure 7 shows the variations of the shadow size $R_{s}$ and the shadow deformation $\delta_{s}$ as functions of the scalar hair parameter $s$ (left column) and the charge parameter $Q$ (right column), for BH spin parameters $a=0.3$ and $a=0.5$. In the top row, $R_{s}$ decreases monotonically with increasing $s$ (left panel, $Q=0.3$ fixed) and with increasing $Q$ (right panel, $s=0.1$ fixed), indicating that both scalar hair and electric charge reduce the apparent size of the shadow. Notably, the curves for $a=0.3$ and $a=0.5$ are nearly indistinguishable in both panels, implying that the variation of $R_{s}$ is insensitive to the spin parameter $a$ within this parameter range. In the bottom row, $\delta_{s}$ increases with both $s$ (left panel, $Q=0.3$) and $Q$ (right panel, $s=0.1$), reflecting growing deviations from circularity as either parameter increases. In contrast to $R_{s}$, the deformation $\delta_{s}$ depends strongly on $a$: for both $s$ and $Q$ scans, the $\delta_{s}$ curve for $a=0.5$ lies systematically above that for $a=0.3$, demonstrating that higher spin enhances the shadow’s distortion. This indicates that while $R_{s}$ is primarily governed by $s$ and $Q$, $\delta_{s}$ is significantly influenced by the BH’s rotation.
V CONSTRAINTS FROM THE EHT OBSERVATIONS OF SUPERMASSIVE BLACK HOLES OF SGR A
The EHT observations of Sgr A impose significant constraints on theoretical models of BHs. By imaging the BH shadow and its surrounding emission, the EHT data enables precise constraints on fundamental parameters. These observations provide unprecedented insights into the nature of BHs and their immediate environments. In this section, we use these data to constrain the parameter space of rotating charged BHs with scalar hair.
@@ V CONSTRAINTS FROM THE EHT OBSERVATIONS OF SUPERMASSIVE BLACK HOLES OF SGR A
Figure 8: Density plots of the shadow angular diameter $\theta_{d}$ for rotating charged black holes with scalar hair at a viewing angle $\theta_{0}=17^{\circ}$. The left panel fixes the charge parameter $Q=0.3$, while the right panel fixes the scalar hair parameter $s=0.1$. The dashed black line indicates $\theta_{d}=46.9\,\mu\mathrm{as}$, corresponding to the lower limit of the angular diameter inferred from EHT observations of Sgr A.
[…]
Figure 9: Density plots of the shadow angular diameter $\theta_{d}$ for rotating charged black holes with scalar hair at a viewing angle $\theta_{0}=90^{\circ}$. The left panel fixes the charge parameter $Q=0.3$, while the right panel fixes the scalar hair parameter $s=0.1$. The dashed black line indicates $\theta_{d}=46.9\,\mu\mathrm{as}$, corresponding to the lower limit of the angular diameter inferred from EHT observations of Sgr A.
@@ VI Conclusion
In this study, we have derived a rotating charged BH solution with scalar hair by applying the NJA to the spherically symmetric charged BH with scalar hair. We have systematically investigated the physical properties of the event horizon and ergosurface in this spacetime. Our analysis shows that the number and structure of horizons depend on the interplay between the charge parameter $Q$, the scalar hair parameter $s$, and the spin parameter $a$. Notably, for fixed $a$ and $Q$, both the radius of the outer horizon and the outer ergosphere decrease as $s$ increases. Similarly, for fixed $a$ and $s$, increasing the charge $Q$ leads to a reduction in both the outer horizon and ergosphere radii.
Additionally, we have applied the Novikov-Thorne model to study thin accretion disks around rotating charged BHs with scalar hair. For fixed spin parameter $a$, the energy flux and radiation temperature of the disk exhibit qualitatively similar radial profiles: they rise with radius, reach a maximum, and then decline. Both quantities increase with either the charge parameter $Q$ or the scalar hair parameter $s$, indicating enhanced radiative efficiency in the presence of larger $Q$ or $s$. Conversely, for fixed $Q$ or $s$, the peak values of the energy flux and temperature decrease as $a$ increases, suggesting that higher spin reduces the peak emission in this class of BH models with scalar hair.
Furthermore, we have analyzed how the parameters $a$, $Q$, and $s$ influence the shape and size of the shadow cast by rotating charged BHs with scalar hair. Our results show that, for fixed spin $a$, the shadow angular diameter decreases as either the charge $Q$ or the scalar hair parameter $s$ increases. In near-extremal cases, the shadow develops a cuspy edge, reflecting strong light bending in the vicinity of the horizon. Comparing with EHT observations of Sgr A, we obtain the constraints $0<Q<0.522745$ (at $s=0.1$) and $0<s<0.283373$ (at $Q=0.3$). These results demonstrate that rotating charged BHs with scalar hair are compatible with current BH imaging data.